OpenStudy (anonymous):

Paulie multiplies two numbers whose product is 1. If one of these numbers is 2, what is the other number?

OpenStudy (lgbasallote):

pauli multiplies two numbers so xy it says the pproduct is 1 so xy = 1 one of the numbers is 2 so let x = 2 2y = 1 what would be y?

OpenStudy (anonymous):

it is 2

OpenStudy (lgbasallote):

not exactly... 2 x 2 = 4 not 1

OpenStudy (anonymous):

is it a one step equation

OpenStudy (lgbasallote):

i believe so

OpenStudy (lgbasallote):

you only do one step from 2y = 1

OpenStudy (anonymous):

i got y=2

OpenStudy (lgbasallote):

what did you do?

OpenStudy (anonymous):

first i did it like this 2y=1 *2=*2 then i eliminated the both of the two's on the left side and then i brought down the y and did 1*2 on the right side and got y=2 that is the way my teacher taught me

OpenStudy (lgbasallote):

you dont multiply by 2...if you multiply both sides by 2 you'll get 4y = 2

OpenStudy (anonymous):

so then what do u do?

OpenStudy (lgbasallote):

you \(\textbf{*divide*}\). remember \[\frac{2}{2} = 1 \; ; \; \frac{3}{3} = 1\; ; \; \frac 44 = 1\; ; \; \text{etc.}\]

OpenStudy (anonymous):

i just worked it out and i got 2